OpenStudy (anonymous):

The length of a rectangle is 5yd more than twice its width, and the area of the rectangle is 63yd^2. Find the dimensions of the rectangle.

OpenStudy (aivantettet26):

Given: Area of the rectangle = 63yd^2 Length = 2W+5 yd Width = W

OpenStudy (aivantettet26):



OpenStudy (aivantettet26):



OpenStudy (aivantettet26):

Area = Length x Width 63 yd^2 = (2w+5) x w distribute w 63 yd^2 =2w^2 +5w equate to zero 2w^2 +5w - 63yd^2 =0 use quadratic equation (w - 4.5) (w + 7) transpose value w = 4.5 w = -7 Since w = -7 is a negative value we wont use it because it almost impossible that length is a negative value. we will then use w =4.5 Since LENGTH = 2W + 5 L = 2(4.5) + 5 L = 14 Therefore we conclude that W = 4.5 yds and L = 14 yds

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Thankkkk you!